In a certain gas system,the pressure remains constant at 150 kpa even though 1000J of heat is added. The volume change is 1.8L. how much does the internal energy in joules of the system change.


1
Expert's answer
2021-10-28T08:52:58-0400

Solution.

"p=150\sdot 10^3 Pa;"

"Q=1000J;"

"\Delta V=1.8L=1.8\sdot 10^{-3} m^3;"

"U -?"

"Q=U+A;"

"A=p\Delta V;"

"A=150\sdot 10^3 Pa\sdot 1.8\sdot 10^{-3} m^3=270J;"

"U=Q-A;"

"U=1000J-270J=730J;"

Answer: "U=730J."